For low-pass RL filter with R = 20 Ω and L = 2 mH , the corner frequency is:


Select one:

a. 3,979 Hz

b. 796 Hz

c. 1,592 Hz

d. 19,894 Hz

e. 7.958 Hz


1
Expert's answer
2021-06-08T09:27:02-0400

"f_c=\frac{R}{2 \pi L}=\frac{20}{2 \pi *2*10^{-3}}=1591.5 Hz"

"\therefore" Correct option is C